# Texture or Semantics? Vision-Language Models Get Lost in Font Recognition

### v1

@@ 6 Categories of Errors
Table 2: Error ratios for four error types across all false predictions on both versions.
@@
7 Do VLMs Truly Focus on Edge Regions?
Different VLMs utilize distinct mechanisms for processing image inputs. Considering both their popularity and open-source availability, we analyze the attention matrices from two models: Llama-3.2-11B-Vision-Instruct and Qwen2-VL-7B-Instruct.
@@ 7 Do VLMs Truly Focus on Edge Regions?
Figure 5: Attention heatmaps from the final cross-attention layer of the Llama-3.2-11B-Vision-Instruct, showing how input text tokens attend to image patches for six different input images from the hard version of the benchmark.
Figure 6: Attention heatmaps from the final layer of the Qwen2-VL-7B-Instruct, depicting the attention distribution of input text tokens over image patches for six different input images from the easy version of the benchmark.
@@ 8 Conclusion
In this paper, we introduce a font recognition benchmark based on 15 popular font styles, designed with both easy and hard versions to comprehensively evaluate the capabilities of current VLMs in this fine-grained task. We conduct extensive experiments with various VLMs and find that they struggle with font recognition, particularly in the hard version of the benchmark, where the stroop effect significantly deteriorates accuracy. To further investigate VLMs’ capabilities in this task, we evaluate different inference strategies, including CoT prompting, MCQ, and few-shot learning. However, our results indicate that these approaches provide minimal improvement in accuracy. Finally, we analyze the attention weights of two open-source VLMs and observe that they struggle to focus on the edge information of characters in the input images, which hinders their ability to accurately recognize font styles and makes them vulnerable to variations in texture content.
@@ 9 Limitation
In this paper, we propose a benchmark for the comprehensive evaluation of VLMs’ performance on font recognition and analyze the factors contributing to their suboptimal performance. However, this study has the following limitations: (i) Due to budget and GPU resource constraints, we are unable to conduct few-shot inference on more powerful models such as Claude-3.5-Sonnet and GPT-4o; future research could further explore these models. (ii) As certain high-performing models are closed-source, we are unable to analyze their attention matrices, limiting our understanding of their decision-making processes. (iii) Our experiments rely on straightforward prompts for inference. Designing more sophisticated prompts, such as guiding models to focus on key regions in the input image, may enhance their performance on this task. We plan to explore this direction in future work.
@@ Acknowledgments
The authors would like to thank the anonymous reviewers for their discussion and feedback.
@@
Ethical considerations play a central role in this research. All models used in this study are either open-source or widely adopted within the scientific community, ensuring transparency and reproducibility. The proposed benchmark and evaluation framework aim to assess the capabilities of current VLMs on font recognition without introducing or reinforcing harmful biases. No personally identifiable information or sensitive data is involved in this work. We are committed to responsible research practices, and we advocate for the transparent reporting and ethical deployment of AI technologies in ways that serve the broader interests of society.
@@ Appendix B Vision-Language Models
Closed-Source Models: We evaluate two multimodal models developed by OpenAI: GPT-4o and GPT-4o-mini, along with two widely recognized models from Anthropic and Google: Claude-3.5-Sonnet and Gemini-2.0-Flash-001.
Open-Source Models: Our experiments include a diverse range of open-source VLMs, spanning nine models across multiple series. Specifically, we test:
@@ B.1 Closed-Source Models
B.2 Open-Source Models
Qwen2-VL-7B-Instruct: Part of the Qwen multimodal series, Qwen2-VL-7B-Instruct is an open-source vision-language model by Alibaba. It provides a balance between efficiency and accuracy for multimodal reasoning tasks.
@@ C.2 Model Implementation
For all VLMs used in the experiments, we set the temperature to 0.0 and top_p to 0.95 to ensure faithfulness and minimize generation randomness, thereby enhancing reproducibility. For open-source models, we obtain the weights from HuggingFace and perform inference using bfloat16 precision. All experiments involving these models are conducted on a setup with four NVIDIA A100 80GB GPUs. For closed-source models, we access the official API for inference. All experiments are conducted between January 25, 2025, and March 25, 2025. To compute the similarity between different images, we use the openai/clip-vit-base-patch32111https://huggingface.co/openai/clip-vit-base-patch32 model on HuggingFace.
@@
Appendix D Capability Gaps Between Open-source and Closed-source Models
@@
The experimental results indicate that contemporary VLMs struggle with font recognition tasks, with a significant accuracy gap persisting between open-source and closed-source models.
@@
In sentence-based font recognition tasks, which is more commonly encountered in real-world scenarios, three out of four closed-source VLMs achieve at least 18% accuracy across various inference settings, with the best-performing model reaching nearly 67% under the multiple-choice question (MCQ) setting. In contrast, among all open-source VLMs evaluated in this study, only Llama-3.2-11B-Vision-Instruct and Qwen2-VL-72B-Instruct attain approximately 18% accuracy, while the remaining models fall below 10%, with some approaching 0% accuracy. These findings highlight the substantial performance disparity between open-source and closed-source VLMs in font recognition.
@@
For the more challenging version of the task, overall performance remains low across both open-source and closed-source models. The hard version introduces the stroop effect, adding an additional cognitive challenge that negatively impacts nearly all VLMs. Under the MCQ setting, all closed-source models are able to achieve accuracy above 6.67%, indicating that, in some cases, they successfully identify font styles rather than being misled by the textual content in the image. In contrast, open-source models fail to do so in almost all cases, with the only exception of Llama-3.2-90B-Vision-Instruct.
Appendix E Related Work
E.1 Vision-Language Models
With the foundation laid by Transformer and CLIP (Vaswani et al., 2023; Radford et al., 2021), Vision-Language Models (VLMs) have emerged as powerful multimodal systems that seamlessly integrate both visual and linguistic modalities. These models bridges the gap between image and text understanding, enabling them to process and interpret complex multimodal data efficiently (OpenAI et al., 2024; Bai et al., 2025).
In recent years, the performance of VLMs has witnessed remarkable improvements, largely driven by the availability of large-scale image-text datasets and the widespread adoption of contrastive learning techniques. These advancements have endowed VLMs with robust general representation capabilities, allowing them to learn meaningful alignments between visual and textual information. Their effectiveness is further demonstrated by their ability to achieve state-of-the-art results across multiple benchmarks, solidifying their role as foundational technologies in multimodal AI research (Zhang et al., 2024b; Bordes et al., 2024).
The strong multimodal capabilities of VLMs make them valuable tools in real-world applications. Models such as BLIP (Li et al., 2022a), Flamingo (Alayrac et al., 2022) and later GPT-4o exhibit impressive performance in visual question answering, image captioning, and multimodal dialogue, enabling applications in interactive AI agents and assistive technologies for visually impaired individuals (Agrawal et al., 2016; Zhang et al., 2024a; Guo et al., 2024; Gao et al., 2024; Zhai et al., 2024). These advancements underscore the versatility of VLMs in handling complex multimodal tasks.
E.2 Font Recognition
Font recognition, a critical task for people such as researchers and designers, inherently combines visual and textual components. However, it has received limited attention within the Vision-Language Models (VLMs) research community. Traditionally, it has been approached as a text-image classification problem, predominantly relying on CNN-based methods (Wang et al., 2015; Tensmeyer et al., 2017; Qi et al., 2025; Wang & Zong, 2023; Mohammadian et al., 2022; Cui & Inoue, 2021). While these approaches achieve relatively high accuracy with their proposed CNN-based architectures on specific datasets, they typically require extensive fine-tuning on large-scale datasets. This requirement makes them time-consuming and also limits their generalization capabilities.
Recently, VLMs have become essential tools in various visual tasks, demonstrating strong adaptability due to their large-scale pretraining data and comprehensive understanding of diverse image types. Given these advantages, VLMs present a promising alternative for font recognition, offering a potentially free and convenient solution. However, there is currently no comprehensive evaluation to systematically assess the capabilities of VLMs in this specific task. This gap in research motivates our study, aiming to provide a thorough evaluation of VLMs for font recognition and to identify their potential limitations.
@@ E.2 Font Recognition
Appendix F Comparative Analysis of Model Accuracy Across Different Fonts
@@
Appendix G Experimental Prompts
We present the simple prompts employed in the experiments shown in Figures 27 through 30.
Appendix H Additional Attention Visualizations
Due to space constraints in the main paper, we give additional visualizations for two open-source models to illustrate their attention patterns on the same input image across different layers. For Llama-3.2-11B-Vision-Instruct, we present attention matrices from all cross-attention layers, highlighting how attention evolves throughout the network. These visualizations reveal that, regardless of whether a layer is shallow or deep, the model consistently struggles to attend to the edge regions of characters in the image, which are critical for accurate font recognition.
@@ Appendix H Additional Attention Visualizations
Zero-Shot Prompt What font is used in the text of the image?
Figure 27: Zero-Shot prompt for all experiments in this paper.
Figure 28: Zero-Shot CoT prompt for all experiments in this paper.
@@
Figure 31: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct, illustrating the progression of attention across the model.
@@
Figure 32: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct, illustrating the progression of attention across the model.
@@
Figure 33: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct, illustrating the progression of attention across the model.
@@
Figure 34: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct, illustrating the progression of attention across the model.
@@
Figure 35: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct, illustrating the progression of attention across the model.
@@
Figure 36: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct, illustrating the progression of attention across the model.
@@
Figure 37: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct, illustrating the progression of attention across the model.
@@
Figure 38: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct, illustrating the progression of attention across the model.
@@
Figure 39: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct, illustrating the progression of attention across the model.
@@
Figure 40: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct, illustrating the progression of attention across the model.
@@
Figure 41: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct, illustrating the progression of attention throughout the model.
@@
Figure 42: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct, illustrating the progression of attention throughout the model.
@@
Figure 43: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct, illustrating the progression of attention throughout the model.
@@
Figure 44: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct, illustrating the progression of attention throughout the model.
@@
Figure 45: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct, illustrating the progression of attention throughout the model.
@@
Figure 46: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct, illustrating the progression of attention throughout the model.
@@
Figure 47: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct, illustrating the progression of attention throughout the model.

### v2

@@ 6 Categories of Errors
Table 2: Error ratios for four distinct error categories, aggregated over all incorrect predictions by 13 vision-language models on both benchmark tiers (easy and hard).
@@
7 Do VLMs Truly Focus on Crucial Edge Regions?
Different VLMs utilize distinct mechanisms for processing image inputs. Considering both their popularity and open-weight availability, we analyze the attention matrices from two models: Llama-3.2-11B-Vision-Instruct and Qwen2-VL-7B-Instruct.
@@ 7 Do VLMs Truly Focus on Crucial Edge Regions?
Figure 5: Attention heatmaps from the final cross-attention layer of Llama-3.2-11B-Vision-Instruct, illustrating how text tokens attend to image patches for six input images from the hard version of the benchmark, all of which the model fails to classify correctly. Attention weights are averaged across all heads.
Figure 6: Attention heatmaps from the middle cross-attention layer ($14^{th}$) of Qwen2-VL-7B-Instruct, illustrating how text tokens attend to image patches for six input images from the easy version of the benchmark, all of which the model fails to classify correctly. Attention weights are averaged across all heads.
@@ 8 Conclusion
In this paper, we introduce a specially designed Font Recognition Benchmark (FRB) comprising 15 popular font styles, designed with both easy and hard versions to comprehensively assess the capabilities of contemporary VLMs in this fine-grained recognition task. We conduct extensive experiments with a range of VLMs and observe that these models consistently struggle with font recognition, particularly on the hard version of the benchmark, where the stroop effect substantially reduces accuracy. This finding highlights that VLMs remain susceptible to misleading visual cues even when performing semantically grounded tasks. To further examine the capacity of VLMs in this setting, we evaluate various inference strategies, including CoT prompting, Multiple-Choice Question answering, and Few-Shot Learning. However, our results demonstrate that these methods offer only marginal improvements in overall accuracy. Additionally, we categorize the error types exhibited by VLMs into four distinct categories and analyze the attention patterns of two publicly available VLMs. Our analysis reveals that these models fail to adequately focus on the edge information of characters in the input images, which impairs their ability to correctly identify font styles and renders them vulnerable to variations in texture content.
@@ 9 Limitation
In this paper, we propose a specially designed benchmark comprising two task versions to enable a comprehensive evaluation of VLMs’ performance on font recognition and analyze the factors contributing to their suboptimal performance. However, our study still has the following limitations: (i) Due to budget constraints, we are unable to conduct few-shot inference on more powerful models such as Claude-3.5-Sonnet and GPT-4o. (ii) As certain high-performing models are closed-source, we are unable to analyze their attention matrices, limiting our understanding of their decision-making processes. (iii) Our experiments rely on straightforward prompts for inference. Designing more sophisticated prompts, such as guiding models to focus on key regions in the input image, may enhance their performance on this task. We plan to explore these directions in future work.
@@ Acknowledgments
The work is partially supported by the NSF of the United States Grant CRII 2451683, an NVIDIA Academic Grants Program, University of California at Merced, and a UC Merced Faculty Research Award. The views and conclusions are those of the authors and should not reflect the official policy or position of the U.S. Government.
@@
Ethical considerations play a central role in this research. All models used in this study are either open-weight or widely adopted within the scientific community, ensuring transparency and reproducibility. The proposed benchmark and evaluation framework aim to assess the capabilities of current VLMs on font recognition without introducing or reinforcing harmful biases. No personally identifiable information or sensitive data is involved in this work. We are committed to responsible research practices, and we advocate for the transparent reporting and ethical deployment of AI technologies in ways that serve the broader interests of society.
@@ Appendix B Vision-Language Models
Closed-Source Models: We evaluate two multimodal models developed by OpenAI: GPT-4o and GPT-4o-mini (OpenAI et al., 2024b; OpenAI, 2024), along with two widely recognized models from Anthropic and Google: Claude-3.5-Sonnet and Gemini-2.0-Flash-001 (Anthropic, 2024; Google, 2025).
Open-Weight Models: Our experiments include a diverse range of open-weight VLMs, spanning nine models across multiple series. Specifically, we test:
@@ B.1 Closed-Source Models
B.2 Open-Weight Models
Qwen2-VL-7B-Instruct: Part of the Qwen multimodal series, Qwen2-VL-7B-Instruct is an open-weight vision-language model by Alibaba. It provides a balance between efficiency and accuracy for multimodal reasoning tasks.
@@ C.2 Model Implementation
For all VLMs used in the experiments, we set the temperature to 0.0 and top_p to 0.95 to ensure faithfulness and minimize generation randomness, thereby enhancing reproducibility.
@@
For open-weight models, we obtain the weights from HuggingFace and perform inference using bfloat16 precision. All experiments involving these models are mainly conducted on a setup with four NVIDIA A100 80GB GPUs.
For closed-source models, we access the official API for inference. All experiments are conducted between January 25, 2025, and March 25, 2025. To compute the similarity between images for selecting few-shot demonstrations, we employ the openai/clip-vit-base-patch32 model on HuggingFace.
@@
Appendix D Capability Gaps Between Open-weight and Closed-source Models
@@
The experimental results indicate that contemporary VLMs struggle with font recognition tasks, with a significant accuracy gap persisting between open-weight and closed-source models.
@@
In the sentence-based font recognition task, which is more commonly encountered in real-world scenarios, three out of four closed-source VLMs achieve at least 18% accuracy across various inference settings, with the best-performing model reaching nearly 67% under the multiple-choice question (MCQ) setting. In contrast, among all open-weight VLMs evaluated in this study, only Llama-3.2-11B-Vision-Instruct and Qwen2-VL-72B-Instruct attain approximately 18% accuracy, while the remaining models fall below 10%, with some approaching 0% accuracy. These findings highlight the substantial performance disparity between open-weight and closed-source VLMs in font recognition.
@@
For the more challenging version of the task, overall performance remains low across both open-weight and closed-source models. The hard version introduces the stroop effect, adding an additional cognitive challenge that negatively impacts nearly all VLMs. Under the MCQ setting, all closed-source models are able to achieve accuracy above 6.67%, indicating that, in some cases, they successfully identify font styles rather than being misled by the textual content in the image. In contrast, open-weight models fail to do so in almost all cases, with the only exception of Llama-3.2-90B-Vision-Instruct.
@@
Appendix E Comparative Analysis of Model Accuracy Across Different Fonts
@@
Appendix F Experimental Prompts
We present the simple prompts employed in the experiments shown in Figures 7 through 10.
Zero-Shot Prompt What font is used in the text of the image?
Figure 7: Zero-Shot prompt for all experiments in this paper.
Figure 8: Zero-Shot CoT prompt for all experiments in this paper.
Appendix G Additional Attention Visualizations
Due to space constraints in the main part of the paper, we give additional visualizations for two open-weight models to illustrate their attention patterns on the same input image across different layers. For Llama-3.2-11B-Vision-Instruct, we present attention matrices from all cross-attention layers, highlighting how attention evolves throughout the network. These visualizations reveal that, regardless of whether a layer is shallow or deep, the model consistently struggles to attend to the edge regions of characters in the image, which are critical for accurate font recognition.
@@ Appendix G Additional Attention Visualizations
Figure 31: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 32: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 33: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 34: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 35: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 36: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 37: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 38: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 39: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 40: Visualization of attention weights from six cross-attention layers of Llama-3.2-11B-Vision-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 41: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 42: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 43: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 44: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 45: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 46: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.
@@
Figure 47: Visualization of attention weights from the 17th layer to the final attention layer of Qwen2-VL-7B-Instruct (averaged across all attention heads), illustrating the progression of attention across the model’s layers.